# Understanding Self-Supervised Learning via Gaussian Mixture Models

### v2

@@ Appendix B Main Proposition
In this section, we state and prove a key proposition which is used to prove our main result (Thm 5.2)
@@
Suppose $F$ parameterized by $\{w_{k},{\bm{\mu}}_{k},I\}_{k\in[K]}$ be a spherical gaussian mixture model (Def LABEL:def:isotropic_gmm) and $\hat{F}$ be its augmentation-enabled gaussian mixture with bias $\delta$ (Def 5.1). Let $S_{F}$ be the fisher subspace (Eqn 1) of $F$ and ${{\bm{A}}}^{*}$ be the optimal solution of the InfoNCE loss (Eqn 2) :
@@ B.1 Column space of A is a subset of Fischer Subspace
$({{\bm{A}}}^{T}{{\bm{x}}})^{T}({{\bm{A}}}^{T}{{\bm{y}}})={{\bm{x}}}^{T}{{\bm{A}}}{{\bm{A}}}^{T}{{\bm{y}}}={{\bm{x}}}^{T}{{\bm{B}}}{{\bm{y}}}$
, where ${{\bm{B}}}={{\bm{A}}}{{\bm{A}}}^{T}$, i.e. ${{\bm{B}}}$ is positive semi-definite (PSD) matrix of rank $r$. We substitute into InfoNCE loss to get :
$\mathcal{L}=-\mathbb{E}_{{{\bm{x}}},\hat{{{\bm{x}}}}}[{{\bm{x}}}^{T}{{\bm{B}}}\hat{{{\bm{x}}}}]+\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}\tilde{{{\bm{x}}}})])]$
@@
$\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}\bar{{{\bm{B}}}}\tilde{{{\bm{x}}}})])]$ $=\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}^{*}\tilde{{{\bm{x}}}})])]$
@@
$=$ $({{\bm{x}}}-{\bm{\mu}}_{z}+{\bm{\mu}}_{z})^{T}{{\bm{R}}}{{\bm{B}}}^{*}{{\bm{R}}}(\tilde{{{\bm{x}}}}-{\bm{\mu}}_{\tilde{z}}+{\bm{\mu}}_{\tilde{z}})$
$=$ $(({{\bm{x}}}-{\bm{\mu}}_{z})^{T}{{\bm{R}}}+{\bm{\mu}}_{z}^{T}{{\bm{R}}}){{\bm{B}}}^{*}({{\bm{R}}}(\tilde{{{\bm{x}}}}-{\bm{\mu}}_{\tilde{z}})+{{\bm{R}}}{\bm{\mu}}_{\tilde{z}})$
$=$ $(({{\bm{x}}}-{\bm{\mu}}_{z})^{T}{{\bm{R}}}+{\bm{\mu}}_{z}^{T}){{\bm{B}}}^{*}({{\bm{R}}}(\tilde{{{\bm{x}}}}-{\bm{\mu}}_{\tilde{z}})+{\bm{\mu}}_{\tilde{z}})$
@@
$\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}^{*}\tilde{{{\bm{x}}}})])]=\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}\bar{{{\bm{B}}}}\tilde{{{\bm{x}}}})])]$
@@ B.2 Column space of A is equal to Fischer Subspace
Without loss of generality assume ${{\bm{v}}}^{T}{{\bm{u}}}_{i}=0\ \forall i\in\mathcal{I}$ (if not then project onto the null space and re-normalize). Now since $\lambda_{j}=0\ \forall i\notin\mathcal{I}$, we can rotate the eigenvectors $[{{\bm{u}}}_{j}]_{j\notin\mathcal{I}}$ with a unitary matrix such that ${{\bm{u}}}_{j}=v$ for some $j$. This operation doesn’t change ${{\bm{B}}}^{*}$
@@ Appendix C Proof for InfoNCE loss
In this section, we use Prop B.1 to prove our main theorem. We recommend the reader to first go through Prop B.1 and it’s proof (in Suppl B) to understand the proof for the main theorem.
@@
$\mathcal{L}=-\mathbb{E}_{{{\bm{x}}},\hat{{{\bm{x}}}}}[{{\bm{x}}}^{T}{{\bm{B}}}\hat{{{\bm{x}}}}]+\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}\tilde{{{\bm{x}}}})])]$
@@
$+\mathbb{E}_{{{\bm{x}}}^{\prime}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}^{\prime}}[\exp({{\bm{x}}}^{\prime T}{\bm{\Sigma}}^{\frac{1}{2}}{{\bm{B}}}{\bm{\Sigma}}^{\frac{1}{2}}\tilde{{{\bm{x}}}}^{\prime})])]$
@@
$\mathcal{L}=-\mathbb{E}_{{{\bm{x}}}^{\prime},\hat{{{\bm{x}}}}^{\prime}}[{{\bm{x}}}^{\prime T}{{\bm{B}}}^{\prime}\hat{{{\bm{x}}}}^{\prime}]+\mathbb{E}_{{{\bm{x}}}^{\prime}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}^{\prime}}[\exp({{\bm{x}}}^{\prime T}{{\bm{B}}}^{\prime}\tilde{{{\bm{x}}}}^{\prime})])]$
@@ Appendix D Proof for SimSiam Loss
In this section, we use the same methodology as in part of Prop B.1 to prove the theorem. First we use the strict convexity of the loss to show that the solution lies in the fisher subspace. Afterwards, by differentiating the loss function w.r.t. any direction in the fisher subspace, we show that all directions in the subspace should be learnt assuming sufficient capacity.
@@
$\mathcal{L}_{SS}({{\bm{A}}})$ $=\underset{({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}}{-\ \mathbb{E}}\Bigg{[}({{\bm{A}}}^{T}{{\bm{x}}})^{T}{{\bm{A}}}^{T}\hat{{{\bm{x}}}}\Bigg{]}+\xi\underset{{{\bm{x}}}\sim F}{\mathbb{E}}\Bigg{[}||{{\bm{A}}}^{T}{{\bm{x}}}||^{2}\Bigg{]}$
@@
[…]
@@
[…]
@@
$\langle{{\bm{B}}}^{*},{{\bm{R}}}^{T}{{\bm{R}}}\rangle=\operatorname{tr}({{\bm{R}}}{{\bm{B}}}^{*}{{\bm{R}}}^{T})=\sum_{i}\lambda_{i}||{{\bm{R}}}{{\bm{u}}}_{i}||_{2}^{2}<\sum_{i}\lambda_{i}||{{\bm{u}}}_{i}||_{2}^{2}=\langle{{\bm{B}}}^{*},{{\bm{I}}}\rangle$
The strict inequality is due to fact that ${{\bm{B}}}^{\prime}{{\bm{v}}}\neq 0$, i.e. there exists a $i$ with $\lambda_{i}>0$ and ${{\bm{v}}}^{T}{{\bm{u}}}_{i}\neq 0$ (and hence $||{{\bm{R}}}{{\bm{u}}}_{i}||_{2}^{2}<||{{\bm{u}}}_{i}||_{2}^{2}$). $||{{\bm{R}}}{{\bm{u}}}_{j}||_{2}^{2}\leq||{{\bm{u}}}_{j}||_{2}^{2}$ is true generally because ${{\bm{R}}}$ is a projection matrix.
@@
if $\xi<\frac{\delta{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}{1+{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}$ for all directions ${{\bm{v}}}$ in $\mathrm{Span}\big{(}{{\bm{M}}}^{\prime}\big{)}$ (as ${{\bm{u}}}_{i}\in\mathrm{Span}\big{(}{{\bm{M}}}^{\prime}\big{)}$), then $\frac{\partial\mathcal{L}_{SS}}{\partial\lambda_{i}}$ is $<0$. $\frac{\delta{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}{1+{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}$ is monotonic in ${{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}$. The minimum value of ${{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}$ is the smallest non-zero eigenvalue of ${{\bm{M}}}^{\prime}$ denoted by $\lambda_{\mathrm{min}}$. Hence if $0<\xi<\frac{\delta\lambda_{\mathrm{min}}}{1+\lambda_{\mathrm{min}}}$ we are good.
@@ Appendix E Proof for CLIP Loss
In this section, we use the first part of Prop B.1 to prove the theorem, i.e. we use the proof of $\mathrm{Col}({{\bm{A}}}^{*})\subseteq S_{F}$.
@@
=-= - $\mathbb{E}_{{{\bm{x}}}_{t},{{\bm{x}}}_{v}}[{{\bm{x}}}^{T}_{t}{{\bm{B}}}{{\bm{x}}}_{v}]+\mathbb{E}_{{{\bm{x}}}_{t}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}_{v}}[\exp({{\bm{x}}}^{T}_{t}{{\bm{B}}}\tilde{{{\bm{x}}}}_{v})])]$
@@ Appendix F Representation collapse in non-contrastive learning
Consider a two component GMM in a $d$ dimensional ambient space. The means of the components lie on the $x$ and $y$ axis (i.e. the first two dimensions), equidistant from the origin. Both components have isotropic covariance. We plot the first two dimensions in the figure below 3.
Figure 3: Caption
Note that for this setup the rank of the fisher subspace is two (i.e., $K=2$) and the fisher subspace is the $x-y$ plane.
@@
But for non-contrastive objectives like SimSiam, while we can only prove that the optimal solution lies in the $x-y$ plane. SimSiam objective is not able to distinguish between the lines $x+y=0$ and $x=y$, and hence might lead to collapse of representations. But as stated in the theorem LABEL:thm:gmm_simsiam, this is only if $r<K$. For $r\geq K$, the SimSiam objective learns the complete fisher subspace (and only the fisher subspace).
Appendix G Experimental Details
G.1 Data Generation
We also ensure that . For directions orthogonal to the mean-subspace we upscale the variance by a factor of $\kappa$. Ambient Dimension We consider the ambient dimension to be 100 for all experiments (except for Scaling plot Fig 2(d), where we let ambient dimension equal to $K-1$). Means of Components For generation means of the components, we first sample $K-1$ times from a 0 mean, ${{\bm{I}}}$ normal distribution in ambient dimension space. We chose the $K$th mean such that the means sum upto 0. Hence the means lie in a $K-1=9$ dimensional space. Covariance Matrix To construct a covariance matrix we start with a unit variance ${{\bm{I}}}$ matrix. We then upscale the variance in (Ambient - NumMeans+1) dimensional space orthogonal to the means subspace by factor of $\kappa$, where $\kappa=\sigma_{\mathrm{max}}({\bm{\Sigma}})/\sigma_{\mathrm{min}}({\bm{\Sigma}})$. We take $\kappa$ to be $1/0.1=10.0$ by default. For scaling experiments, we choose random NumMeans/2 orthogonal directions, and increase the covariance in those directions. Note for condition number experiment we take $\kappa$ to be $1/0.x$ where $x$ goes from 1 to 9.
G.2 Orthonormalized InfoNCE and SimSiam
We plot the ARI/AMI numbers for orthonormalized InfoNCE and SimSiam matrices in Fig 4. Specifically, we do a QR decomposition on these matrices and take the orthogonal Q matrix as the projection matrix. In these plots the convergence issues of InfoNCE loss become more apparent, as we can see that InfoNCE lagging behind optimal.
@@ G.2 Orthonormalized InfoNCE and SimSiam
Figure 4: We further extend the results presented in Fig 2 with orthonormalized InfoNCE and SimSiam mappings

### v3

@@ Appendix B Main Proposition
In this section, we state and prove a key proposition which is used to prove our main result (Thm 4.1)
@@
Suppose $F$ parameterized by $\{w_{k},{\bm{\mu}}_{k},I\}_{k\in[K]}$ be a spherical gaussian mixture model and $\hat{F}$ be its augmentation-enabled gaussian mixture with bias $\delta$ (Def 2.2). Let $S_{F}$ be the fisher subspace (Eqn 1) of $F$ and $\bm{A}$$^{*}$ be the optimal solution of the InfoNCE loss (Eqn 2) :
@@ B.1 Column space of A is a subset of Fischer Subspace
$({{\bm{A}}}^{T}{{\bm{x}}})^{T}({{\bm{A}}}^{T}{{\bm{y}}})={{\bm{x}}}^{T}{{\bm{A}}}{{\bm{A}}}^{T}{{\bm{y}}}={{\bm{x}}}^{T}{{\bm{B}}}{{\bm{y}}},$
where ${{\bm{B}}}={{\bm{A}}}{{\bm{A}}}^{T}$, i.e. $\bm{B}$}} is positive semi-definite (PSD) matrix of rank $r$. We substitute into InfoNCE loss to get :
$)]\displaystyle\mathcal{L}=-\mathbb{E}_{{{\bm{x}}},\hat{{{\bm{x}}}}}[{{\bm{x}}}^{T}{{\bm{B}}}\hat{{{\bm{x}}}}]+\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}\tilde{{{\bm{x}}}}$)])]
@@
$)]\displaystyle\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}\bar{{{\bm{B}}}}\tilde{{{\bm{x}}}}$)])] $)]\displaystyle=\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}^{*}\tilde{{{\bm{x}}}}$)])]
@@
[…]
@@
$)]\displaystyle\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}^{*}\tilde{{{\bm{x}}}})])]=\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}\bar{{{\bm{B}}}}\tilde{{{\bm{x}}}}$)])]
@@ B.2 Column space of A is equal to Fischer Subspace
Without loss of generality assume ${{\bm{v}}}^{T}{{\bm{u}}}_{i}=0\ \forall i\in\mathcal{I}$ (if not then project onto the null space and re-normalize). Now since $\lambda_{j}=0\ \forall i\notin\mathcal{I}$, we can rotate the eigenvectors ${{\bm{u}}}_{j}]_{j\notin\mathcal{I}}$ with a unitary matrix such that ${{\bm{u}}}_{j}=v$ for some $j$. This operation doesn’t change $\bm{B}$$^{*}$
@@ Appendix C Proof for InfoNCE loss
In this section, we use Proposition˜B.1 to prove our main theorem. We recommend the reader to first go through Proposition˜B.1 and it’s proof (in Appendix˜B) to understand the proof for the main theorem.
@@
$)]\displaystyle\mathcal{L}=-\mathbb{E}_{{{\bm{x}}},\hat{{{\bm{x}}}}}[{{\bm{x}}}^{T}{{\bm{B}}}\hat{{{\bm{x}}}}]+\mathbb{E}_{{{\bm{x}}}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}}[\exp({{\bm{x}}}^{T}{{\bm{B}}}\tilde{{{\bm{x}}}}$)])]
@@
$)]\displaystyle+\mathbb{E}_{{{\bm{x}}}^{\prime}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}^{\prime}}[\exp({{\bm{x}}}^{\prime T}{\bm{\Sigma}}^{\frac{1}{2}}{{\bm{B}}}{\bm{\Sigma}}^{\frac{1}{2}}\tilde{{{\bm{x}}}}^{\prime}$)])]
@@
$)]\displaystyle\mathcal{L}=-\mathbb{E}_{{{\bm{x}}}^{\prime},\hat{{{\bm{x}}}}^{\prime}}[{{\bm{x}}}^{\prime T}{{\bm{B}}}^{\prime}\hat{{{\bm{x}}}}^{\prime}]+\mathbb{E}_{{{\bm{x}}}^{\prime}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}^{\prime}}[\exp({{\bm{x}}}^{\prime T}{{\bm{B}}}^{\prime}\tilde{{{\bm{x}}}}^{\prime}$)])]
@@ Appendix D Proof for SimSiam Loss
In this section, we use the same methodology as in part of Proposition˜B.1 to prove the theorem. First we use the strict convexity of the loss to show that the solution lies in the fisher subspace. Afterwards, by differentiating the loss function w.r.t. any direction in the fisher subspace, we show that all directions in the subspace should be learnt assuming sufficient capacity.
@@
[…]
@@
$=-\langle{{\bm{A}}}{{\bm{A}}}^{T},\delta\underset{k}{\sum}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}\rangle+\xi\langle{{\bm{A}}}{{\bm{A}}}^{T},\underset{k}{\sum}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}+{\bm{\Sigma}}\rangle$
@@
where we have ${{\bm{B}}}={{\bm{A}}}{{\bm{A}}}^{T}$, ${{\bm{M}}}=\underset{k}{\sum}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}$, ${{\bm{B}}}^{\prime}={\bm{\Sigma}}^{\frac{1}{2}}{{\bm{B}}}{\bm{\Sigma}}^{\frac{1}{2}}$ and ${{\bm{M}}}^{\prime}={\bm{\Sigma}}^{-\frac{1}{2}}{{\bm{M}}}{\bm{\Sigma}}^{-\frac{1}{2}}$. Note that the loss is linear in $\bm{B}$}} (and $\bm{B}$$^{\prime}$). Since we restrict ourselves to ${{\bm{A}}}||_{2}\leq 1$, it is also true that ${{\bm{B}}}||_{2}\leq 1$.
@@
[…]
@@
if $\xi<\frac{\delta{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}{1+{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}$ for all directions $\bm{v}$}} in $\mathrm{Span}\big({{\bm{M}}}^{\prime}\big)$ (as ${{\bm{u}}}_{i}\in\mathrm{Span}\big({{\bm{M}}}^{\prime}\big)$), then $\frac{\partial\mathcal{L}_{SS}}{\partial\lambda_{i}}$ is <0<0. $\frac{\delta{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}{1+{{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}}$ is monotonic in ${{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}$. The minimum value of ${{\bm{v}}}^{T}{{\bm{M}}}^{\prime}{{\bm{v}}}$ is the smallest non-zero eigenvalue of $\bm{M}$$^{\prime}$ denoted by $\lambda_{\mathrm{min}}$. Hence if $0<\xi<\frac{\delta\lambda_{\mathrm{min}}}{1+\lambda_{\mathrm{min}}}$ we are good.
@@ Appendix E Proof for CLIP Loss
In this section, we use the first part of Proposition˜B.1 to prove the theorem, i.e. we use the proof of $\mathrm{Col}({{\bm{A}}}^{*})\subseteq S_{F}$.
@@
$=$- $\mathbb{E}_{{{\bm{x}}}_{t},{{\bm{x}}}_{v}}[{{\bm{x}}}^{T}_{t}{{\bm{B}}}{{\bm{x}}}_{v}]+\mathbb{E}_{{{\bm{x}}}_{t}}[\log(\mathbb{E}_{\tilde{{{\bm{x}}}}_{v}}[\exp({{\bm{x}}}^{T}_{t}{{\bm{B}}}\tilde{{{\bm{x}}}}_{v})])]$
@@ Appendix F Representation collapse in non-contrastive learning
Consider a two component GMM in a $d$ dimensional ambient space. The means of the components lie on the $x$ and $y$ axis (i.e. the first two dimensions), equidistant from the origin. Both components have isotropic covariance. We plot the first two dimensions below in Figure˜3.
Figure 3: The means of the GMM components in the first two dimensions.
Note that for this setup the rank of the fisher subspace is two (i.e., $K=2$) and the fisher subspace is the xyx-y plane.
@@
But for non-contrastive objectives like SimSiam, while we can only prove that the optimal solution lies in the xyx-y plane. SimSiam objective is not able to distinguish between the lines $x+y=0$ and $x=y$, and hence might lead to collapse of representations. But as stated in the Theorem˜4.1, this is only the case if $r<K$. For $r\geq K$, the SimSiam objective learns the complete Fisher subspace (and only the fisher subspace).
Appendix G Experimental Details and Additional Results
@@
Figure 4: We further extend the results presented in Figure˜2 with orthonormalized InfoNCE and SimSiam mappings
G.1 Data Generation for Synthetic Experiments
Ambient Dimension We consider the ambient dimension to be 100 for all experiments (except for scaling curves in Figure˜2(d), where we let ambient dimension equal to K1K-1).
Means of Components For generating means of the components, we first sample K1K-1 times from a 0 mean, $\bm{I}$}} normal distribution in ambient dimension space. We chose the $K$th mean such that the means sum upto 0. Hence the means lie in a K1=9K-1=9 dimensional space.
Covariance Matrix To construct a covariance matrix we start with a unit variance $\bm{I}$}} matrix. We then upscale the variance in (Ambient - NumMeans+1) dimensional space orthogonal to the means subspace by factor of $\kappa$, where $\kappa=\sigma_{\mathrm{max}}({\bm{\Sigma}})/\sigma_{\mathrm{min}}({\bm{\Sigma}})$. We take $\kappa$ to be $1/0.1=10.0$ by default. For scaling experiments, we choose random NumMeans/2 orthogonal directions, and increase the covariance in those directions. Note for condition number experiment we take $\kappa$ to be $1/0.x$ where $x$ goes from 1 to 9.
G.2 Orthonormalized InfoNCE and SimSiam
We plot the ARI/AMI numbers for orthonormalized InfoNCE and SimSiam matrices in Figure˜4. Specifically, we do a QR decomposition on these matrices and take the orthogonal Q matrix as the projection matrix. In these plots the convergence issues of InfoNCE loss become more apparent, as we can see that InfoNCE lagging behind optimal.
G.3 CIFAR-100 Experiments and Additional Results
In this section, we present a seeded version of our results in the manuscript for the CIFAR-100 experiments. We randomly sample 20 classes from the datasets and collect the data points that belong to these 20 classes. The randomly selected subset of images are then normalized before inputting the the mapping method. Once the methods learn linear maps, it is then applied on the normalized data. Finally, we cluster the (linearly) transformed data using K-Means where the number of classes is the same as the number of components in the GMM. We repeat this procedure for 5 times with different seeds for the random number generator. We report the results in Table˜3, which are consistent with the results in the main text.
Table 3: Clustering performance of linear dimensionality reduction methods on CIFAR100: Bolded and underlined values indicate the best and second-best scores for each column. We report clustering performance for K-Means on $r=5,10,15,19$-dimensional subspaces (mapping dimensions) using the linear mappings learned by 6 different methods. We measure the clustering performance using ARI and AMI. We run the methods for 5 different seeds, corresponding to 5 random 20-class subsets of CIFAR-100 datasets. The average scores over 5 runs are reported with standard deviation given in paranthesis. InfoNCE and LDA outperform the rest by significant margins in all mapping dimensions. LDA shows the best performance in terms of ARI, while InfoNCE achieves higher AMI values across the board.
Mapping Dim = 5 Mapping Dim = 10 Mapping Dim = 15 Mapping Dim = 19
Method ARI AMI ARI AMI ARI AMI ARI AMI
Random 0.020 (0.003) 0.061 (0.009) 0.026 (0.003) 0.078 (0.009) 0.028 (0.003) 0.084 (0.007) 0.033 (0.007) 0.094 (0.015)
Ambient 0.039 (0.006) 0.113 (0.013) 0.039 (0.006) 0.113 (0.013) 0.039 (0.006) 0.113 (0.013) 0.039 (0.006) 0.113 (0.013)
PCA 0.038 (0.006) 0.110 (0.013) 0.040 (0.006) 0.113 (0.014) 0.040 (0.006) 0.114 (0.014) 0.041 (0.007) 0.116 (0.015)
SimSiam 0.035 (0.009) 0.111 (0.020) 0.037 (0.009) 0.113 (0.021) 0.035 (0.010) 0.110 (0.021) 0.035 (0.009) 0.110 (0.018)
InfoNCE 0.058 (0.009) 0.146 (0.016) 0.064 (0.011) 0.156 (0.018) 0.063 (0.005) 0.157 (0.014) 0.066 (0.008) 0.159 (0.014)
LDA 0.060 (0.008) 0.145 (0.014) 0.074 (0.012) 0.153 (0.013) 0.074 (0.010) 0.148 (0.014) 0.076 (0.014) 0.149 (0.017)